# How many points converge to a heteroclinic network in an aperiodic way?

### v1

@@ 4. Aperiodic sequences
If $\mathbf{q}$ is f.a.s., then $\Sigma^{k}\mathbf{q}$ is f.a.s. by inclusion of the respective sets. Since $\mathbf{q}$ is aperiodic, there is a countably infinite set of sequences $\Sigma^{k}\mathbf{q}$ which are distinct and all f.a.s., proving the assertion. ∎
@@
As a direct consequence the asymptotic basins of attraction of two sequences that are not related by $\Sigma$ are also disjoint. With this we can state our next result that limits the total number of f.a.s. sequences that may coexist in a given network.
@@
:=$\mathcal{F}:=\{\mathcal{S}_{\delta}(\mathbf{q})\mid\mathbf{q}\in\mathbf{Q}\text{ and }\mu(\mathcal{S}_{\delta}(\mathbf{q}))>0\},$
@@ 5.1. Geometric restrictions on the realized sequences
The possibility for the realization of aperiodic sequences near the Kirk–Silber network [19] is restricted by the fact that it is impossible for trajectories near the network to go from one cycle to the other and back. In the terminology of Example 2.4 this means that $\mathrm{ABA}$ and $\mathrm{BAB}$ cannot appear in any realized sequence. In fact, results in [27] imply that the only sequences realized in the Kirk–Silber system are of the form $\mathrm{A}^{k}\mathrm{BBB}\dotsb$ with arbitrary $k\in\mathbb{N}$ (or with the roles of $\mathrm{A}$ and $\mathrm{B}$ reversed, depending on the choice of parameters). This fully characterizes the set of sequences that are realized in the Kirk–Silber system. In particular, all realized sequences are preperiodic and thus no aperiodic sequence is realized (much less f.a.s.) here.
@@
Podvigina [15] discusses a network in $\mathbb{R}^{6}$ where these restrictions do not apply. Its graph can be created from the Kirk–Silber network by adding two new nodes, thus replacing both $[\xi_{3}\to\xi_{1}]$ and $[\xi_{4}\to\xi_{1}]$ by $[\xi_{3}\to\xi_{3}^{\prime}\to\xi_{1}]$ and $[\xi_{4}\to\xi_{4}^{\prime}\to\xi_{1}]$, respectively [^3] We use a different labeling for the nodes than Podvigina to highlight the similarity to the Kirk–Silber network.; see Figure 3. By slight abuse of terminology we label the cycles $[\xi_{1}\to\xi_{2}\to\xi_{3}\to\xi_{3}^{\prime}\to\xi_{1}]$ and $[\xi_{1}\to\xi_{2}\to\xi_{4}\to\xi_{4}^{\prime}\to\xi_{1}]$ as $\mathrm{A}$ and $\mathrm{B}$, respectively. Podvigina shows that under suitable assumptions the periodic sequence (which she calls an omnicycle) defined through $\overline{\mathrm{A}\mathrm{B}}$ is f.a.s. (which she calls trail-stable).
More generally, as pointed out in Section 2, results in [20] show that having only real eigenvalues prevents infinite switching and restricts the set of finite sequences that may be realized. However, this does not necessarily mean that no aperiodic sequences can be realized (and potentially be f.a.s.) in such systems. On the one hand, Postlethwaite and Dawes [14] derived restrictions for possible periodic ways to approach a heteroclinic network with strong stability properties. While they observe irregular cycling dynamics towards the network, the stability of particular aperiodic sequences remains unclear. On the other hand, Podvigina presents numerical evidence hinting at potentially stable aperiodic dynamics in [15], but at the same time states that, numerically, aperiodic sequences are practically indistinguishable from long periodic ones.
[…]
@@
The authors of [7] offer similar results for a network in the context of the Rock-Paper-Scissors-Lizard-Spock game: They prove fragmentary asymptotic stability for a number of periodic sequences, but also encounter regions in parameter space where they cannot find stable periodic behavior. Their numerical results seem to indicate stable aperiodic dynamics, but a proof of aperiodicity, in the sense that there is an aperiodic f.a.s. sequence, is lacking.
@@ 5.3. Two homoclinic loops
In our third example, we discuss a different potential source for the emergence of f.a.s. aperiodic sequences in a product system of two stable homoclinic cycles with distinct convergence speeds. Specifically, we discuss a toy example of two homoclinic loops to derive heuristic arguments for the presence or absence of f.a.s. aperiodic sequences in a certain class of systems. Consider a product system in $\mathbb{R}^{2}\times\mathbb{R}^{2}$ with a hyperbolic saddle equilibrium at the origin and two attracting homoclinic trajectories (one in each factor of the system) converging to the origin in forward and backward time. This produces a homoclinic network N comprised of two cycles, say $\mathrm{A}$ and $\mathrm{B}$. Typically, this network is attracting in such a product system (rather than the product of the cycles) [29]. This means trajectories approaching this attractor can be assigned a sequence in these two symbols corresponding to successive transitions along one cycle or the other. In contrast to [30], where the absense of finite switching was proved for such a system, we here focus on the possibility to obtain aperiodic sequences as points approach N.
@@
which implies that the relation we were looking for is (asymptotically) linear. Despite the simplifications this can be seen as an indication that in the setting of homoclinic cycles with exponential stability one should not be surprised to encounter preperiodic rather than aperiodic sequences for typical trajectories.
@@ 6. Discussion
Third, what happens under time-reversal? Asymptotic stability relates to convergence as time increases. Thus, it is natural to ask what the asymptotic dynamics are as time is reversed and whether it relates to the network itself. A trajectory through $x\in W^{\textrm{u}}(\xi_{p})\cap\mathcal{A}_{\delta}(\mathbf{q})$ approaches N in forward time (along $\mathbf{q}$) and in backward time (to $\xi_{p}$); this relates to heteroclinic connections of higher depth, see, for example, [31, 32]. More generally, given an aperiodic f.a.s. sequence $\mathbf{q}$, which (finite) sequences can we put in front of $\mathbf{q}$ such that the resulting sequence is also f.a.s.? Finally, can we have a meta-homoclinic (or bi-directional) heteroclinic sequence in the sense that there are trajectories converging to a network in forward and backward time, and if so, are they f.a.s.?

### v2

@@ 4. Aperiodic sequences
If $\mathbf{q}$ is f.a.s., then by definition $\mu(\mathcal{B}_{\delta}(\mathbf{q}))>0$. Since $\mathcal{B}_{\delta}(\mathbf{q})\subset\mathcal{B}_{\delta}(\Sigma\mathbf{q})$, we have $\mu(\mathcal{B}_{\delta}(\Sigma\mathbf{q}))\geq\mu(\mathcal{B}_{\delta}(\mathbf{q}))>0$. In the same way it follows that
$\mu(\mathcal{B}_{\delta}(\Sigma^{k}\mathbf{q}))\geq\mu(\mathcal{B}_{\delta}(\Sigma^{k-1}\mathbf{q}))\geq\ldots\geq\mu(\mathcal{B}_{\delta}(\mathbf{q}))>0.$
Thus, $\Sigma^{k}\mathbf{q}$ is f.a.s. for any $k\in\mathbb{N}$. Since $\mathbf{q}$ is aperiodic, there is a countably infinite set of sequences $\Sigma^{k}\mathbf{q}$ which are distinct and all f.a.s., proving the assertion. ∎
@@
As a direct consequence the asymptotic basins of attraction of two sequences that are not related by $\Sigma$ are also disjoint. With this we can state our next result that—through measure-theoretic arguments—limits the total number of f.a.s. sequences that may coexist in a given network.
@@
$\mathcal{F}:=\{\mathcal{S}_{\delta}(\mathbf{q})\mid\mathbf{q}\in\mathbf{Q}\text{ and }\mu(\mathcal{S}_{\delta}(\mathbf{q}))>0\},$
@@ 5.1. Geometric restrictions on the realized sequences
Near the Kirk–Silber network [19] aperiodic sequences are not realized because it is impossible for trajectories within a small neighborhood to go from one cycle to the other and back; cf. Example 2.10. If the symbols $\mathrm{A}$ and $\mathrm{B}$ correspond to the constituent cycles (as in Example 2.5), this means that $\mathrm{ABA}$ and $\mathrm{BAB}$ cannot appear in any realized sequence. In fact, results in [27] imply that the only sequences realized in the Kirk–Silber system are of the form $\mathrm{A}^{k}\mathrm{BBB}\dotsb$ with arbitrary $k\in\mathbb{N}$ (or with the roles of $\mathrm{A}$ and $\mathrm{B}$ reversed, depending on the choice of parameters). This fully characterizes the set of sequences that are realized in the Kirk–Silber system. In particular, all realized sequences are preperiodic and thus no aperiodic sequence is realized (much less f.a.s.) here.
@@
Podvigina [15] discusses a network in $\mathbb{R}^{6}$ where these restrictions do not apply. Its graph can be created from the Kirk–Silber network by adding two new nodes, thus replacing both $\xi_{3}\to\xi_{1}]$ and $\xi_{4}\to\xi_{1}]$ by $\xi_{3}\to\xi_{3}^{\prime}\to\xi_{1}]$ and $\xi_{4}\to\xi_{4}^{\prime}\to\xi_{1}]$, respectively [^3] We use a different labeling for the nodes than Podvigina to highlight the similarity to the Kirk–Silber network.; see Figure 3. By slight abuse of terminology we label the cycles $\xi_{1}\to\xi_{2}\to\xi_{3}\to\xi_{3}^{\prime}\to\xi_{1}]$ and $\xi_{1}\to\xi_{2}\to\xi_{4}\to\xi_{4}^{\prime}\to\xi_{1}]$ as $\mathrm{A}$ and $\mathrm{B}$, respectively. Podvigina shows that under suitable assumptions the periodic sequence (an omnicycle) defined through $\overline{\mathrm{A}\mathrm{B}}$ is f.a.s. (which she calls trail-stable).
More generally, as pointed out in Section 2, results in [20] show that having only real eigenvalues prevents infinite switching and restricts the set of finite sequences that may be realized. However, this does not necessarily mean that no aperiodic sequences can be realized (and potentially be f.a.s.) in such systems. On the one hand, Postlethwaite and Dawes [14] derived restrictions for possible periodic ways to approach a heteroclinic network with strong stability properties. While they observe irregular cycling dynamics in the sense that all periodic sequences on the network they consider are unstable, the stability of individual aperiodic sequences remains unclear. On the other hand, Podvigina presents numerical evidence hinting at potentially stable aperiodic dynamics in [15], but at the same time states that, numerically, aperiodic sequences are practically indistinguishable from long periodic ones.
$\xi_{1}$$\xi_{2}$$\xi_{3}$$\xi_{3}^{\prime}$$\xi_{4}$$\xi_{4}^{\prime}$$\mathrm{A}$$\mathrm{B}$
@@
The authors of [7] offer similar results for a network in the context of the Rock-Paper-Scissors-Lizard-Spock game: They prove fragmentary asymptotic stability for a number of periodic sequences, but also encounter regions in parameter space where they cannot find stable periodic behavior. Their numerical results indicate stable aperiodic dynamics, but a proof of aperiodicity, in the sense that there is an aperiodic f.a.s. sequence, is lacking.
@@ 5.3. Two homoclinic loops
In our third example, we discuss a different potential source for the emergence of f.a.s. aperiodic sequences in a product system of two stable homoclinic cycles with distinct convergence speeds. Specifically, we study a toy example of two homoclinic loops to derive heuristic arguments for the presence or absence of f.a.s. aperiodic sequences in a certain class of systems. The main insight is that if the equilibrium is hyperbolic, one should expect preperiodic rather than aperiodic sequences. This indicates possible obstructions for the emergence of f.a.s. aperiodic sequences and suggests future work with a focus on two homoclinic loops from a nonhyperbolic equilibrium.
Consider a product system in $\mathbb{R}^{2}\times\mathbb{R}^{2}$ with a hyperbolic saddle equilibrium at the origin and two attracting homoclinic trajectories (one in each factor of the system) converging to the origin in forward and backward time. This produces a homoclinic network N comprised of two cycles, say $\mathrm{A}$ and $\mathrm{B}$. Typically, this network is attracting in such a product system (rather than the product of the cycles) [29]. This means trajectories approaching this attractor can be assigned a sequence in these two symbols corresponding to successive transitions along one cycle or the other. In contrast to [30], where the absence of finite switching was proved for such a system, we here focus on the possibility to obtain aperiodic sequences as points approach N.
@@
which implies that the relation we were looking for is (asymptotically) linear. Despite the simplifications, this indicates that sequences close to the homoclinic loops are predominantly preperiodic rather than aperiodic.
This suggests that—in the context of a system with two homoclinic loops—hyperbolicity of the equilibrium can be an obstruction to the emergence of f.a.s. aperiodic sequences. At the same time, it suggests that f.a.s. aperiodic sequences may appear as a bifurcating case when the equilibrium is nonhyperbolic. We discuss this further below.
@@ 6. Discussion
Third, what happens under time-reversal? Asymptotic stability relates to convergence as time increases. Thus, it is natural to ask what the asymptotic dynamics are as time is reversed and whether it relates to the network itself. A trajectory through $x\in W^{\textrm{u}}(\xi_{p})\cap\mathcal{A}_{\delta}(\mathbf{q})$ approaches N in forward time (along $\mathbf{q}$) and in backward time (to $\xi_{p}$); this relates to heteroclinic connections of higher depth; see, for example, [31, 32]. More generally, given an aperiodic f.a.s. sequence $\mathbf{q}$, which (finite) sequences can we put in front of $\mathbf{q}$ such that the resulting sequence is also f.a.s.? Finally, can we have a bi-directional heteroclinic sequence in the sense that there are trajectories converging to a network in forward and backward time, and if so, are they f.a.s.?
Acknowledgements. We are grateful to P. Ashwin, S. Castro, C. Postlethwaite, A. Rodrigues, and A. Rucklidge for helpful discussions on the content of this paper. We also thank an anonymous reviewer for their comments and questions that helped improve the exposition.